# Cut-based Conflict Analysis in Mixed Integer Programming

### v1

@@ 9             $C_{\textnormal{confl}}\leftarrow C_{\textnormal{reason}}$
In Example 3 we show that no globally valid inequality can be constructed that is violated in the local bounds by using only the reason constraint $C_{2}$, the conflict constraint $C_{3}$ and the global bounds of the variables. Next we show that, after resolving all non-relaxable continuous variables from the reason constraint $C_{2}$, the new reason constraint contains only binary variables and propagates $x_{1}$:
@@
This constraint contains only binary variables, propagates $x_{1}$ to 1 non-tightly, and hence can be reduced as shown in Section 4.
@@ 7 Conflict Analysis for Mixed Integer Programs
Next, we briefly discuss the limitations of the conflict analysis algorithm under the presence of general integer variables. Again, the goal is to resolve a bound change on a variable by some suitable reduction and aggregation of constraints. As expressed in Proposition 1, whenever the propagation of the resolved variable is tight, we can simply aggregate the reason for the bound change to the current conflict and the resolvent remains infeasible. However, in the presence of general integer variables, our reduction algorithms are not guaranteed to find a reduced reason that propagates tightly.
Non-tight propagation occurs because of the existence of non-integer vertices for the linear relaxation of $C_{\textnormal{reason}}$ over the local domain $L$. The goal then is to find a cut that separates the non-integer vertex from the feasible region $C_{\textnormal{reason}}\cap G$, where $G$ is the box defined by the global bounds. This is not possible if the general integer variable contributes with a non-global bound to the propagation of the reason constraint. Figure 2 gives an example that illustrates such a situation.
Figure 2: Example where it is impossible to find a linear cut that separates the non-integer vertex (2,0.5) from the feasible region (green points).
One approach to addressing the resolution of general integer variables is proposed in Jovanovic and de Moura (2013). Here the authors allow only fixings of variables to their current lower or upper bound. While this strategy restricts the branching heuristic, it also avoids scenarios like the one illustrated in Figure 2, allowing reduction algorithms to effectively eliminate non-integer vertices.
@@ 8 Experimental Evaluation
All techniques discussed in this paper have been implemented in the open-source MIP solver SCIP 9.0 Bolusani et al. (2024). The experiments were conducted on a cluster equipped with Intel Xeon Gold 5122 CPUs @ 3.60GHz and 96GB of RAM. Each run was performed on a single thread with a time limit of two hours. To mitigate the effects of performance variability (Lodi and Tramontani 2013) and to ensure a fair comparison of the different conflict analysis algorithms, we used a large and diverse set of test instances, namely the MIPLIB 2017 benchmark set (Gleixner et al. 2021), with five permutations of each individual model. After excluding all models that are not solvable by any setting for any of the five seeds, we obtained a total of 795 measurements per run. For the remainder of this paper, we will refer to the combination of a model and a permutation as an instance.
@@ 8.1 Implementation details
In our experiments, we do not use the Chvátal-Gomory-based reduction since it is dominated by the cMIR-CA reduction (and does not generalize to problems with continuous variables). Moreover, as demonstrated in Proposition 4, opting for complementation over weakening results in more robust reasoning constraints. Therefore, we consider the cMIR-CA reduction solely with complementation.
@@ 8.2 Comparison of Conflict Analysis Algorithms in Isolation
Table 1 summarizes the results of the experiment for the three different conflict analysis algorithms. The table shows the number of instances (inst), the number of instances solved to optimality (opt), the average solving time in seconds (time(s)), the average number of conflicts generated (confs), the average number of non-zeros in conflicts (avg.length), the percentage of conflicts that were used in propagation (used(%)), and the average number of bound changes applied (per node) by the conflicts (bdchgs).
@@
Two of the most noticeable differences in Table 1 are the large discrepancy in the number of propagations and the percentage of conflicts actually used in propagation. Both are significantly larger for the cut-based variants, which is a favorable result. When analyzing our results, we found three different reasons for this behavior.
@@ 8.3 MIP Performance
In this experiment, we activate cut-based conflict analysis as an additional conflict analysis method within SCIP. This means that conflicts from both, cut- and graph-based conflict analysis, are generated at each infeasibility. Note that in SCIP, graph-based conflict analysis generates All-FUIP conflicts by default, so-called reconvergence cuts Achterberg (2007a) and also locally valid conflicts, while our cut-based conflict analysis implementation only generates globally valid conflicts and is restricted to 1-UIP conflicts, as a result of preliminary experiments. These differences, together with the significantly different structure of the generated constraints, serve as a motivation that the two approaches will complement each other.
@@
Table 2: Performance comparison of SCIP vs SCIP with cMIR-CA .
@@
Table 2 shows the performance of SCIP with and without the cMIR-CA on MIPLIB2017. The table is divided into different subsets of instances. The first row shows the performance on all 795 instances, and the second row on all 492 instances that are “affected” by conflict analysis, i.e., where the execution path of the two settings differs.
The next four rows show the performance on instances where at least one of the two settings needs at least 1, 10, 100, or 1000 seconds to solve the instance, respectively. The last two rows show the performance on instances that are solved to optimality by both settings and the subset of those instances that are affected by conflict analysis “aff.-both-solved”. We included these rows because they allow for a comparison of tree sizes.
The results show that using cMIR-CA leads to a notable improvement in the performance of SCIP. Namely, it increases the number of solved instances from 637 to 644 and decreases the overall running time by 5%. On affected instances, we observed a runtime decrease of 8%. As shown in the rows [1,tilim], [10,tilim], [100,tilim], and [1000,tilim], the performance improvement is more pronounced on harder instances, with a speedup of up to 15% on instances that need at least 1000 seconds to solve.
@@ 9 Conclusion
Our path to generalize cut-based conflict analysis to mixed integer problems in full generality was paved by both positive and negative results. First, we showed that the cut-based approach cannot be directly applied to mixed binary programs without considering additional problem information. To address this limitation, we designed a new reduction algorithm that not only utilizes two constraints (reason and conflict) in each iteration of conflict analysis but also incorporates constraints whose propagation is responsible for tightened bounds of continuous variables. We demonstrate that in the presence of general integer variables conflict analysis cannot be guaranteed to work as long as we refrain from lifting the problem into higher dimension. However, our empirical results show that these cases occur only rarely in practice, rendering our cut-based conflict analysis a fruitful approach also for general integer variables.
@@
Acknowledgments
The work for this article has been conducted within the Research Campus Modal funded by the German Federal Ministry of Education and Research (BMBF grant numbers 05M14ZAM, 05M20ZBM). Jakob Nordström is supported by the Swedish Research Council grant 2016-00782 and the Independent Research Fund Denmark grant 9040-00389B
Part of this work was carried out while some of the authors participated in the extended reunion for the program Satisfiability: Theory, Practice, and Beyond at the Simons Institute for the Theory of Computing at UC Berkeley in the spring of 2023. This work has also benefited greatly from discussions during the Dagstuhl Seminar 22411 Theory and Practice of SAT and Combinatorial Solving.

### v2

@@ 9    $C_{\textnormal{confl}}\leftarrow C_{\textnormal{reason}}$
In Example˜2 we show that no globally valid inequality can be constructed that is violated in the local bounds by using only the reason constraint $C_{2}$, the conflict constraint $C_{3}$ and the global bounds of the variables. Next we show that, after resolving all non-relaxable continuous variables from the reason constraint $C_{2}$, the new reason constraint contains only binary variables and propagates $x_{1}$:
@@
This constraint contains only binary variables, propagates $x_{1}$ to 1 non-tightly, and hence can be reduced as shown in Section˜4.
@@ 7 Conflict Analysis for Mixed Integer Programs
Next, we briefly discuss the limitations of the conflict analysis algorithm under the presence of general integer variables. Again, the goal is to resolve a bound change on a variable by some suitable reduction and aggregation of constraints. As expressed in Proposition˜1, whenever the propagation of the resolved variable is tight, we can simply aggregate the reason for the bound change to the current conflict and the resolvent remains infeasible. However, in the presence of general integer variables, our reduction algorithms are not guaranteed to find a reduced reason that propagates tightly.
Non-tight propagation occurs because of the existence of non-integer vertices for the linear relaxation of $C_{\textnormal{reason}}$ over the local domain $L$. The goal then is to find a cut that separates the non-integer vertex from the feasible region $C_{\textnormal{reason}}\cap G$, where $G$ is the box defined by the global bounds. This is not possible if the general integer variable contributes with a non-global bound to the propagation of the reason constraint. Figure˜3 gives an example that illustrates such a situation.
Figure 3: Example where it is impossible to find a linear cut that separates the non-integer vertex (2,0.5) from the feasible region.
One approach to addressing the resolution of general integer variables is proposed in Jovanovic and de Moura (2013). Here the authors allow only fixings of variables to their current lower or upper bound. While this strategy restricts the branching heuristic, it also avoids scenarios like the one illustrated in Figure˜3, allowing reduction algorithms to effectively eliminate non-integer vertices.
@@ 8 Experimental Evaluation
All techniques discussed in this paper have been implemented in the open-source MIP solver SCIP 9.0 Bolusani et al. (2024). We have uploaded the code and results to a public GitHub repository hosted by the INFORMS Journal on Computing Mexi et al. (2024). The experiments were conducted on a cluster equipped with Intel Xeon Gold 5122 CPUs @ 3.60GHz and 96GB of RAM. Each run was performed on a single thread with a time limit of two hours. To mitigate the effects of performance variability (Lodi and Tramontani 2013) and to ensure a fair comparison of the different conflict analysis algorithms, we used a large and diverse set of test instances, namely the MIPLIB 2017 benchmark set (Gleixner et al. 2021), with five permutations of each individual model. After excluding all models that are not solvable by any setting for any of the five seeds, we obtained a total of 795 measurements per run. For the remainder of this paper, we will refer to the combination of a model and a permutation as an instance.
@@ 8.1 Implementation details
In our experiments, we do not use the Chvátal-Gomory-based reduction since it is dominated by the cMIR-CA reduction (and does not generalize to problems with continuous variables). Moreover, as demonstrated in Proposition˜3, opting for complementation over weakening results in more robust reasoning constraints. Therefore, we consider the cMIR-CA reduction solely with complementation.
@@ 8.2 Comparison of Conflict Analysis Algorithms in Isolation
Table˜1 summarizes the results of the experiment for the three different conflict analysis algorithms. The table shows the number of instances (inst), the number of instances solved to optimality (opt), the average solving time in seconds (time(s)), the average number of conflicts generated (confs), the average number of non-zeros in conflicts (avg.length), the percentage of conflicts that were used in propagation (used(%)), and the average number of bound changes applied (per node) by the conflicts (bdchgs).
@@
Two of the most noticeable differences in Table˜1 are the large discrepancy in the number of propagations and the percentage of conflicts actually used in propagation. Both are significantly larger for the cut-based variants, which is a favorable result. When analyzing our results, we found three different reasons for this behavior.
@@ 8.3 MIP Performance
In this experiment, we activate cut-based conflict analysis as an additional conflict analysis method within SCIP. This means that conflicts from both cut- and graph-based conflict analysis, are generated at each infeasibility. For the cut-based conflict analysis, we only use the cMIR-based reduction. As seen in the previous section cMIR-CA and CoefT-CA perform similarly, however in an extended experiment with a longer time limit of 6 hours and 10 different random seeds, the cMIR-based reduction was on average 8% faster on “affected” instances and required 8% fewer nodes on those solved to optimality.
@@
Table 2: Performance comparison of SCIP vs SCIP +cMIR-CA .
@@
Table˜2 shows the performance of SCIP with and without the cMIR-CA on MIPLIB2017. One instance was excluded due to numerical issues, leaving 794 instances in total. The table is divided into different subsets of instances. The first row shows the performance on all 794 instances, and the second row on all 492 instances that are “affected” by conflict analysis, i.e., where the execution path of the two settings differs. The next three rows show the performance on instances where at least one of the two settings needs at least 10, 100, or 1000 seconds to solve the instance, respectively. The next row “diff-timeouts” shows the performance on instances where only one of the two settings is able to solve the instance to optimality. SCIP solved 21 instances that SCIP + cMIR-CA did not, while SCIP + cMIR-CA solved 29 instances that SCIP did not. In a post-hoc analysis with a time limit of 6 hours, there is only a single instance that SCIP solved in under 2 hours that SCIP + cMIR-CA still could not solve after 6 hours. Notably, there are 9 instances that SCIP failed to solve within 6 hours, while SCIP + cMIR-CA had already solved them in under 2 hours. The last two rows show the performance on instances that are solved to optimality by both settings and the subset of those instances that are affected by conflict analysis “aff.-both-solved”. We included these rows because they allow for a comparison of tree sizes.
The results show that using cMIR-CA leads to a notable improvement in the performance of SCIP. Namely, it increases the number of solved instances from 636 to 644 and decreases the overall running time by 5%. On affected instances, we observed a runtime decrease of 8%. As shown in the rows [10,tilim], [100,tilim], and [1000,tilim], the performance improvement is more pronounced on harder instances, with a speedup of up to 15% on instances that need at least 1000 seconds to solve.
@@ 9 Conclusion
On our path to generalize cut-based conflict analysis to mixed integer problems in full generality we encountered both positive and negative results. First, we showed that the cut-based approach cannot be directly applied to mixed binary programs without considering additional problem information. To address this limitation, we designed a new reduction algorithm that not only utilizes two constraints (reason and conflict) in each iteration of conflict analysis but also incorporates constraints whose propagation is responsible for tightened bounds of continuous variables. We demonstrate that in the presence of general integer variables conflict analysis cannot be guaranteed to work as long as we refrain from lifting the problem into higher dimension. However, our empirical results show that these cases occur only rarely in practice, rendering our cut-based conflict analysis a fruitful approach also for general integer variables.